OpenStudy (anonymous):

There are two factors of -36 such that one factor is 11 less than half of the other factor. Choose all the pairs of these factors. -2 and 18 -6 and 6 3 and -12 4 and -9

OpenStudy (radar):

let x and y be the factors of -36 but x must equal (y/2) - 11 Now use algebra, you have two equations and two unknowns.

OpenStudy (radar):

Using substitution: x =-36/y -36/y = (y/2) - 11